OpenStudy (anonymous):

Use implicit differentiation to find the slope of the tangent line to the curve 4xy^3+3xy=14 at the point (2,1) I got 1/30.. can someone check my answer? :)

OpenStudy (anonymous):

4y^3 + 12xy^2y' + 3y + 3xy' = 0 (12xy^2+3x)y'=-4y^3-3y y'=(-y^3-3y)/(12xy^2+3x) (2,1) => (x,y) y' = (-1-3)/(12(2)(1)+3(2)) = -4/30 y' = -2/15

OpenStudy (anonymous):

it says that's wrong.

OpenStudy (anonymous):

y'=(-y^3-3y)/(12xy^2+3x) this part is copied down wrong. should be y'=(-4y^3-3y)/(12xy^2+3x) If you follow my work, you'll notice that I dropped the 4 before the y^3. This changes the answer to -7/30

OpenStudy (anonymous):

ohhh i get! I had my negative and plus signs messed up when I did it.. Thanks!

OpenStudy (anonymous):

happens all the time. After years of calculus, i still make the mistake of missing a coefficient in a previous equation

OpenStudy (anonymous):

I had -(3y+4y^3)/(12xy^2+3x)